system by substitution
x+3y=13
3x+5y=19
i can't seem to get this.....

Answer by MOD(6)

Hi, I'm here to help.
For this problem lets we need to use substition. So we need to solve for x or y in one of the above equations and then substitue that answer for x or y in the other equation.
Start by solving for x in this equation.
x + 3y = 13
Subract 3y from both sides of the equation.
x = 13 - 3y
Now that we have an answer for x we can substitute it as x for the other equation.
3(13 - 3y) + 5y = 19
First we will distribute the 3 to both 13 and -3y.
39 - 9y + 5y = 19
Combine like terms.
39 - 4y = 19
Subtract 39 from both sides of the equation to isolate the y.
-4y = -20
Isolate the y further by dividing both sides of the equation by -4.
y = 5
Now we can substitute y in the first equation to find x.
x + 3(5) = 13
x + 15 = 13
Isolate the x by subtracting 15 from both sides of the equation.
x = -2
So we have that x = -2 and y = 5 as our answers.
Now check your answers.
-2 + 3(5) = 13
-2 + 15 = 13
13 = 13
3(-2) + 5(5) = 19
-6 + 25 = 19
19 = 19